sum of(1+2+4+8+16+.........30th term)
Answer by aswathytony(47)

this is a G.P.( Geometric progression)
sum of n terms of a G.P = a( r^n -1 )/r -1
1 + 2 + 4 + 8 +16 +....+30th term
first term ,a = 1
common ratio, r = 2
number of terms, n =30
substituting in formula, sum of 30terms = 1 ( 2^30 - 1) / 2-1
= 2^30 -1 /1
= 2^30 - 1